OpenStudy (anonymous):

the radius of a wheel is 80 centimeters, if the wheel rotates thru an angle of 60 degrees, how many centimeters does it move? express your answer in terms of pi.

OpenStudy (anonymous):

\[\pi 2*80*\frac{60{}^{\circ}}{360{}^{\circ}}=\frac{80 \pi }{3}=83.7758\text{ cm} \]\[\text{one} \text{ full} \text{ rotation} =\pi 2*80=502.655\text{ cm} \]